OpenStudy (vane11):

Can someone try this and tell me if they got the same plz?? I got height:1, length/width:3 but I feel that's too "pretty" to be right If an open box is made from a tin sheet 6 in. square by cutting out identical squares from each corner and bending up the resulting flaps, determine the dimensions of the largest box that can be made. (Round your answers to two decimal places.)

jimthompson5910 (jim_thompson5910):

What work do you have so far? Can you post that?

OpenStudy (vane11):

(6-2x)(6-2x) x(36-12x-12x+4x^2) x(36-24x+4x^2) 0=36-24x^2+4x^3 and from here I tried both the dy/dx and -b+/-sqrtb-4ac / 2a and got 3 and 1 is my math wrong and I just haven't caught it?? I'll keep looking it over

jimthompson5910 (jim_thompson5910):

you went from x(36-24x+4x^2) to 36-24x^2+4x^3 which is incorrect

jimthompson5910 (jim_thompson5910):

it should be 36x-24x^2+4x^3

jimthompson5910 (jim_thompson5910):

y = 36x-24x^2+4x^3 dy/dx = 36 - 48x + 12x^2 0 = 36 - 48x + 12x^2 Use the quadratic formula to get x = 1, x = 3 You would then use the first derivative test to find out that x = 1 is a local max, x = 3 is a local min

jimthompson5910 (jim_thompson5910):

So you got the height of x = 1 correct The length/width are incorrect though

jimthompson5910 (jim_thompson5910):

since 6-2x = 6-2*1 = 4

OpenStudy (vane11):

ohhhh ok wow I missed the easy part, and just assumed the two values (min/max) where the answers, thank you!

jimthompson5910 (jim_thompson5910):

you're welcome